OpenStudy (ttp):

What is a rule for the function identified by this set of ordered pairs? {(ordered pair one comma one), (ordered pair two comma four), (ordered pair three comma nine), (ordered pair four comma sixteen), (ordered pair five comma twenty-five)} double the number square the number triple the number multiply the number by 4

OpenStudy (anonymous):

square the number

OpenStudy (anonymous):

1^2=1 2^2=4 3^2=9 4^2=16

OpenStudy (ttp):

Thanks !

OpenStudy (anonymous):

hope i helped :)

OpenStudy (anonymous):

welcome!

OpenStudy (ttp):

You did :)